Reports 1-1 of 1 Clear search Modify search
AdV-DET (Commissioning)
mwas, gouaty, magazzu, bulten, tacca, fafone, lumaca, nardecchia, rocchi - 0:39 Thursday 14 January 2021 (50440) Print this report
Summary of the DET shift - OMC tuning and SDB2 alignment

This afternoon, the DET shift was performed in the following conditions : PR, BS, NI aligned, SR in parking position, NE, WI, WE misaligned.

The following activities were performed :

  • At the beginning of the shift, the angular position of the SDB2 bench was adjusted in order to have the B5 beam and B1p beam roughly centered on the cameras. The TY setpoint has been updated in the configuration file TY = +1100. In this condition, the B5 beam is at X=-0.12 / Y=-0.8 on the camera and B1p at X=+0.16 and Y=0.
  • Then the correct propagation of the Hartmann beam was checked as reported in entry 50434.
  • The offsets of the B1p quadrant DC channels were adjusted. The B1p beam was centered on the quadrants, by acting with picomotors on the mirrors B1p_M1/M2. It was then possible to close the centering loop based on the galvos (Fig.1).
  • Similarly, the offsets of the B5 quadrants were corrected. Then the B5 beam was centered on the quadrants, by acting on the mirrors B5_M1/M2. And the galvo loops of the B5 quadrants could be closed (Fig.2).
  • The OMC was locked with B1_PD3 (on the S polarized TEM00) in order to perform a coarse tuning of the OMC polarization matching. To this purpose we acted on the two waveplates controlled from SDB1_Rot Channel 2, axis 2 and Channel 3, axis 1.
  • Then the OMC was locked on a side band TEM00, in order to be able to align the B1_PD1/2 photodiodes without saturating them.
  • Both B1_PD1 and B1_PD2 photodiodes have been aligned with picomotors of the mirror mounts B1_M2 and B1_M3. During this step, we noticed that the shutter of the B1_PD1 photodiode was closed, despite the fact that we had sent a command to open it and the shutter status was marked as open in the DAQ box process. By sending another close and open commands, the shutter effectively opened, and now seems to be working correctly.
  • Then the remaining of the shift was dedicated to the OMC polarization tuning, that we tried to improve by minimizing the power reflected by the OMC in the P polarization at the crossing of the S polarization resonance. The final state of the polarisation tuning is shown in Fig.3 and Fig.4. There was about 0.54 mW of power on B1_PD3 at the resonance of the S polarized TEM00, and the P polarized peak is at 0.012 mW. This corresponds to a polarization mismatch lower than 2.2%.
Images attached to this report
Comments to this report:
mwas - 15:07 Thursday 14 January 2021 (50444) Print this report

In the OMC scan there are symmetrical peaks around the carrier TEM00 polarization S mode. I expect these to be 6MHz and 8MHz lower and upper sideband. The P polarization carrier TEM00 is probably the shoulder on the 6MHz LSB, so the tuning is much better than the 2% of light lost due to polarization mistuning. But it is hard to judge with the 6MHz LSB overlapping.

The expected transmission of the 6MHz is 0.5% due to cavity finesse and length for the S-polarization. But here p-polarization of the 6MHz will be closely overlapping with carrier S-polarization TEM00. If there is no separation we want to have the polarization tuning much better than 0.5%, the overlap is not exact, so the cavity pole will filter it at least by a factor 2. So this is acceptable if the polarization is mistuned by 0.1%. This seems feasible, but probably it is safer to reduce the force applied by the PZT screw on the OMC, to reduce the birefrigence and bring the P-polarization peak more in between the S-polarization peak and the 6MHz LSB.

But in LAPP back in 2015 we have seen this peak depend on the temperature of the OMC, so this could be tested before the intervention on SDB1 end of January. In preparation of that I have changed the OMC temperature set point from 23 to 24 degrees, so that the OMC has the time to thermalize.



Images attached to this comment
Search Help